The stopping distance of a car after the brakes have been applied varies directly as the square of the speed.
A.If a car traveling 60mph can stop in 200ft.,how far will it take a car to stop if it is traveling at 50mph
B.Suppose the car was able to stop in 50ft. At what rate was its traveling?

Answer by Boreal(15235)

d=k*s^2
200=k*60^2
200/3600=k
k=1/18 or 0.057
50=k*s^2
50/(1/18)=s^2
900=s^2
s=30 mph
This makes sense, because as the speed doubles (30 to 60) the stopping distance should increase by a factor of 4 (50 to 200).